# On the Closed-Form of Flow Matching: Generalization Does Not Arise from Target Stochasticity

### v1

@@ C.3 Figure 4
In Figure 4, for each dataset (CIFAR-$10$ and CelebA $64$), one network is trained using a vanilla conditional flow matching with the default parameters of Tong et al. [2024] (the most important ones are recalled in Table 1). Then images are generated first following the closed-form formula of the optimal velocity field $\hat{u}^{\star}$ from 00 to $\tau$. And then following the velocity field learned with a usual conditional flow matching $u_{\theta}$ from $\tau$ to $1$.
C.4 Figure 5
@@
For experiments involving training on CIFAR-10 (Figures 3 and 4), we rely on the standard 34 million parameters U-Net for diffusion by Nichol and Dhariwal [2021], with default settings from the torchfm codebase [Tong et al., 2024]. For each algorithm, the networks are trained for $500$k iterations with batch size $128$, i.e., $1280$ epochs.
@@
For CelebA $64\times 64$ (Figure 4), we rely on the training script of pnpflow library [^8] https://github.com/annegnx/PnP-Flow [Martin et al., 2025], which uses a U-Net from Huang et al. [2021], Ho et al. [2020b].

### v2

@@ D.4 Figure˜3
In Figure˜3, for each dataset (CIFAR-10 and CelebA $64\times 64$), one network is trained using a vanilla conditional flow matching with the default parameters of tong2023improving (the most important ones are recalled in Table˜3). Then images are generated first following the closed-form formula of the optimal velocity field $\hat{u}^{\star}$ from 0 to $\tau$. And then following the velocity field learned with a usual conditional flow matching $u_{\theta}$ from $\tau$ to $1$.
D.5 Figure˜4
@@
For experiments involving training on CIFAR-10 (Figures˜2 and 3), we rely on the standard 34 million parameters U-Net for diffusion by nichol2021improved, with default settings from the torchfm codebase [tong2023improving]. For each algorithm, the networks are trained for 500500k iterations with batch size $128$, i.e., $1280$ epochs.
@@
For CelebA $64\times 64$ (Figure˜3), we rely on the training script of pnpflow library [^4] https://github.com/annegnx/PnP-Flow [martin2024pnpflow], which uses a U-Net from huang2021variational, Ho2020.